OpenStudy (anonymous):

19 - 3(2x - 1)=7 please help me solve this

OpenStudy (anonymous):

19 - 3(2x - 1)=7 -3(2x-1)=7-19 -3(2x-1)=-12 2x-1=4 2x=5 x=5/2

OpenStudy (anonymous):

Solving brackets 19 - 3(2x - 1)=7 Brackets cannot be solved. doing multiply or divide. 19 - 6x -(-3)=7 19 - 6x +3=7 adding or subtracting what ever needs to be done. 22-6x=7 adding 6x on both sides 22-6x+6x=7+6x 22=7+6x subtracting 7 on both sides 22-7=7+6x-7 15=6x 6x=15 divide both sides by 6 \[\frac{ 6x }{ 6 } = \frac{ 15 }{ 6}\] \[x = \frac{ 15 }{ 6 }\] \[x= \frac{ 5 }{ 2 }\] Cant write in improper fraction. converting into decimals x=2.5